# The Future of Large Language Model Pre-training is Federated

### v1

@@ 2.3 Federated Fine-tuning and Parameter Efficient Fine-tuning of LLMs
Until now, full federated pre-trained LLMs have not been accomplished because researchers could not solve the dual challenges of its communication overhead and pre-training large models on resource-challenged devices. That said, researchers, eager to reap the benefits of federated learning, have nonetheless made progress on federating downstream LLM training tasks whose computational and communication demands are lower, such as fine-tuning, parameter-efficient fine-tuning (PEFT), and prompt-tuning.
@@
For example, Hilmkil et al. [53] use FL to fine-tune all parameters of ALBERT [54] and BERT [55], reaching 90% of the accuracy achieved by a centrally trained model on text classification tasks. Meanwhile, Riedel et al. [56] found that BERT fine-tuned in an FL scenario could perform as well as a centralized model on multilingual text classification. Wang et al. [57] and Weller et al. [58] also conducted federated fine-tuning on private data but preceded it with centralized pre-training on public data, finding that this workflow improved fine-tuning accuracy, even when the pre-training was only being done on a 1% sampling of a large public corpus or if the client-side data was non-IID, respectively.
@@
Much progress has also been made on federated PEFT, whose computational and communication hurdles are lower than those of federated fine-tuning. Researchers have shown that a model that has been subject to federated PEFT can outperform the original pre-trained model [59], outperform siloed client models [60], and even outperform federated fine-tuning [61, 62], including in non-IID scenarios [63], but with far lower computation and communication costs because clients only need to update and transmit the smaller set of parameters. Federated LoRA, for example, may consume as little as 0.058% of the communication cost of federated full fine-tuning [60]. To reduce these costs even further, Xu et al. [64] add Noise Contrastive Estimation to reduce memory demands andXu et al. [65]add backpropagation-free training to improve memory and time efficiency. Meanwhile, to address the greater impact that non-IID client distributions can have on federated PEFT performance, Babakniya et al. [66] precede federated LoRA with a round of federated efficient sparse fine-tuning, reducing the performance gap while keeping combined training time low. Differently, Kim et al. [67] abandon a global model altogether, instead using FL to train client-specific models that benefit from sharing data via FL but are more resilient to client drift.
@@
Researchers have also demonstrated that federated prompt tuning, wherein clients tune a set of continuous soft prompts that are appended to input prompts, can perform as well as the PTM [61], the siloed client models [60] and even federated full fine-tuning [68, 69] while lowering communication costs. Accompanied by an adaptive optimization method, it may also reduce the impact of client drift due to non-IID data [70, 71]. P-tuning, a variety of prompt tuning that concatenates discrete text prompts with the continuous prompt embeddings, was also shown to perform as well as the PTM [61] or local transfer tuning [71] on certain tasks, while significantly reducing the number of model parameters that clients must train and transmit.
@@ 3 Design Principles for Federated Generative Pre-Training of LLMs
This section proposes a series of inclusive principles necessary for such federations to be effective and incorporate them into our system design. The principles we chose are meant to tackle the foundational issues of federated LLM pre-training with a particular focus on data and hardware inclusivity, robustness, and efficiency.
Broad Access to Data:
The ability to train an LLM should depend on the data that an actor or set of actors possess rather than unrestricted access to hardware. Thus, we aim to transform organizations that right now can serve as data providers only [11, 72], when their ownership of the data is respected, into active participants in the LLM training process. We believe that incorporating such actors directly into the training process and offering them an incentive to participate, obtaining a model performing well on their data, is the natural next step in the proliferation of generative AI generally and LLMs in particular.
@@
Training should be possible without the strong synchronization requirements of standard data-parallel training [38] to accommodate geographically distributed and poorly connected participants. Our federated learning solutions allow orders-of-magnitude reductions in communication frequency compared to centralized solutions. While such improvements provide an efficiency boost to all actors in the federation, they also provide particular benefits to data sources that may have been underrepresented in the past. For example, incorporating private data silos from regions with a lower internet presence may help alleviate the challenge of NLP for low-resource languages [73, 74].
@@
Organizations with valuable data must be able to participate in the federated network, even with limited hardware resources. We define the minimum set of hardware resources based on the memory requirements of the model; i.e., their GPUs must have sufficient VRAM to hold the model and at least one sample. If a client suffers from either computational or communication limitations, we account for it by adjusting the maximum round duration or allowing them to modulate the amount of local training they undertake. Thus, our system can accommodate a broad set of potential participants with hardware varying from a few medium-powered GPUs to small data centers.

### v2

@@ 2.3 Federated Fine-tuning and Parameter Efficient Fine-tuning of LLMs
Until now, full federated pre-trained LLMs have not been accomplished because researchers could not solve the dual challenges of its communication overhead and pre-training large models on resource-challenged devices. That said, researchers, eager to reap the benefits of federated learning, have nonetheless made progress on federating downstream LLM training tasks whose computational and communication demands are lower, such as fine-tuning, parameter-efficient fine-tuning (PEFT), and prompt-tuning. We outline these developments here for a complete reference.
@@
For example, Hilmkil et al. [53] use FL to fine-tune all the model parameters of ALBERT [54] and BERT [55], reaching 90% of the accuracy achieved by a centrally trained model on text classification tasks. Meanwhile, Riedel et al. [56] found that BERT fine-tuned in an FL setting could perform as well as a centralized model on multilingual text classification tasks. Wang et al. [57] and Weller et al. [58] also conducted federated fine-tuning on private data but preceded it with centralized pre-training on public data. They found that this workflow improved fine-tuning accuracy, even when pre-training only on a 1% sampling of a large public corpus or if the client’s data was non-IID.
@@
Much progress has also been made on federated PEFT, whose computational and communication hurdles are lower than those of federated fine-tuning. Researchers have shown that a model that has been subject to federated PEFT can outperform the original pre-trained model [59], outperform siloed client models [60], and even outperform federated fine-tuning [61, 62], including in non-IID scenarios [63], but with far lower computation and communication costs because clients only need to update and transmit the smaller set of parameters. Federated LoRA, for example, may consume as little as 0.058% of the communication cost of federated full fine-tuning [60]. To reduce these costs even further, Xu et al. [64] add Noise Contrastive Estimation to reduce memory demands and Xu et al. [65] add backpropagation-free training to improve memory and time efficiency. Meanwhile, to address the more significant impact that non-IID client distributions can have on federated PEFT performance, Babakniya et al. [66] precede federated LoRA with a round of federated efficient sparse fine-tuning, reducing the performance gap while keeping combined training time low. Differently, Kim et al. [67] abandon a global model altogether, instead using FL to train client-specific models that benefit from sharing data via FL but are more resilient to client drift.
@@
Federated prompt tuning, wherein clients tune a set of continuous soft prompts appended to input prompts, has also demonstrated its effectiveness. This technique can significantly reduce the number of model parameters that clients must train and transmit compared with full fine-tuning. Models fine-tuned this way perform better than siloed client transfer learning [61], better than centralized training by benefiting from improved generalisation [68, 69], and also reduce the impact of client drift due to non-IID data [70, 71] by leveraging an adaptive optimization method.
@@ 3 Design Principles for Federated Generative Pre-Training of LLMs
This section proposes a series of inclusive principles necessary for such federations to be effective and incorporate them into the design of our system, Photon. The principles we chose are meant to tackle the foundational issues of federated LLM pre-training with a particular focus on data and hardware inclusivity, robustness, and efficiency.
Broad Access to Data and Compute:
The ability to train an LLM should depend on the data that a participant or a group possesses rather than unrestricted access to hardware. Thus, we aim to transform organizations that right now can serve as data providers only [12, 72], when their ownership of the data is respected, into active participants in the LLM training process. We believe that incorporating such contributors directly into the federated learning process and offering them an incentive to participate, obtaining a model performing well on their data, is the natural next step in the proliferation of generative AI generally and LLMs in particular. While some data-rich organizations may be unable or unwilling to invest in computing power, we believe that voluntarily partnering with a compute-rich yet data-poor one in a federated training context provides an excellent avenue for their participation. Thus, Photon aims to create a broad network of data producers and consumers where a single client may offer computing resources, data sources, or both to the training process.
@@
Pre-training should be possible without the strong synchronization requirements of standard data-parallel training [38] to accommodate geographically distributed and poorly connected participants. Our federated learning solutions allow orders-of-magnitude reductions in communication frequency compared to centralized solutions. While such improvements boost efficiency for all participants in the federation, they also offer particular benefits to data sources that may have been underrepresented in the past. For example, incorporating private data silos from regions with a lower internet presence may help alleviate the challenge of NLP for low-resource languages [73, 74].
@@
Organizations with valuable data must be able to participate in the federated network, even with poorly connected or limited hardware resources. In the case of clients that may hold distributed compute nodes lacking the connectivity necessary to support the high-bandwidth Ring AllReduce algorithm necessary for classical data-parallel training (e.g., Infiniband), we perform local federated training over their compute nodes and transparently aggregate the node updates before sending results to the main server. Suppose a client suffers from either computational or communication limitations. In that case, we account for it by adjusting the maximum round duration or allowing them to modulate the amount of local training they undertake. Thus, our system can accommodate a broad set of potential participants with hardware varying from a few medium-powered GPUs to small data centers.